OpenStudy (anonymous):

Help !! Giving Out Medals .. :) Evaluate 2b2 - 3b + 4a2 for a= -6 & b=8 A. -40 B. 202 C. 248 D. 346

OpenStudy (triciaal):

straight substitution rewrite the original where you have a put -6 and when you have b put 8

OpenStudy (triciaal):

2b2 - 3b + 4a2 2(8)^2 -3(8) + 4 (-6)^2 2*64 -24 + 4* 36 4( 32 -6 + 36) = 4*62 = 248 = C